Reduce answer to the lowest terms 5/6+3/4
Answer by Alan3354(30993)

Reduce answer to the lowest terms 5/6+3/4
-----------
Find a common denominator.
The product always works = 6*4 = 24
---
20/24 + 18/24 = 38/24
= 19/12
=================
You hear about LCD, Least Common Denominator, but it's necessary to use that. The product always works. You might save time by not finding the LCD on a test. You just have to reduce the result.
--------------
If the problem says to find and use the LCD, tho, you have to do that.